OpenStudy (darkworld131):

What is the area of the rectangle who's length is 1 1/3 ft and width is 2/3 ft. A: 160in (squared) B:56in (squared) C:28 in (squared) D:80 in (Squared) None of the above

OpenStudy (anonymous):

area = length x width = 1 1/3 x 2/3

OpenStudy (anonymous):

Welcome to O.S. :-D Step 1: Convert your units to inches! Step 2: Area = length * width

OpenStudy (anonymous):

1 + 1/3 = 4/3 (4/3 ft) * \(\large \frac{12 inches}{1 foot}\) = ____ inches (2/3 ft) * \(\large \frac{12 inches}{1 foot}\) = ____ inches ( ____ inches)( ____ inches) = ( _____ inches\(^2\))

OpenStudy (anonymous):

@darkworld131 Make sense now?

OpenStudy (darkworld131):

yes kind of, thanks